OpenStudy (anonymous):

0.285=log(x) ...solve for x.

Vocaloid (vocaloid):

rewrite both sides to base 10

Vocaloid (vocaloid):

10^0.285 = x

OpenStudy (danjs):

\[\log _{b}(x) = b^x\]

OpenStudy (anonymous):

Oh wow awesome thanks!

OpenStudy (danjs):

logs are just exponents

OpenStudy (danjs):

log base b of x = b to what power is x